1. Mike’s Sport Shop deposits $3,600 at the end of each year for 12 years at 7% annual interest.
a. How much will this ordinary annuity be worth at the end of the 12 years? (5 points)
Answer:
b. How much more will this annuity be worth (annuity due) if Mike deposits the money at the beginning of each year instead of at the end of each year? (5 points)
Answer:
2. Barb and John Reed want to know how much they must deposit in a retirement savings account today to have payments of $1,750 every six months for 15 years. The retirement account is paying 8% annual interest, compounded semiannually. (5 points)
Answer:
3. Lena Dimock is saving for her college expenses. She sets aside $200 at the beginning of each three months in an account paying 8% annual interest, compounded quarterly. How much will Lena have accumulated in the account at the end of four years? (5 points)
Answer:

Part II. Case Study
Julie has just completed the rigorous process of becoming a Certified Financial Planner (CFP). She is looking forward to working with individuals on saving for retirement. She would like to show her clients the value of an annuity program as one of the best options for investing current earnings in a tax-deferred account.
1. If a client puts the equivalent of $55 per month, or $660 per year, into an ordinary annuity, how much money would accumulate in 20 years at 3% compounded annually? (5 points)
Answer:

2. Jackie, a 25 year old client, want to retire by age 65 with $2,000,000. How much would she have to invest annually, assuming a 6% rate of return? (5 points)
Answer:

3. Another client, Wynona, decides that she will invest $5,000 per year in a 6% annuity for the first ten years, then $6,000 for the next ten years, and then $4,000 per year for the last ten years, how much will she accumulate? [Hint: Treat each ten-year period as as separate annuity and compute the Future Value. After the ten years, assume that the value will continue to grow at compound interest for the remaining years of the 30 years. Use tables from Unit 6 to compute compound interest.

Found 2 solutions by edward d, Hyot:
Answer by edward d(1)

Lena Dimock is saving for her college expenses. She sets aside $200 at the beginning of each three months in an account paying 8% annual interest, compounded quarterly. How much will Lena have accumulated in the account at the end of four years

Answer by Hyot(1)

3. Lena Dimock is saving for her college expenses. She sets aside $200 at the beginning of each three months in an account paying 8% annual interest, compounded quarterly. How much will Lena have accumulated in the account at the end of four years? (5 points)


Answer: Number of years times number of intrest periods per year = 4(4) = 16
Annual intrest rate/Number of periods per year = 8/4 = 2%
$200 ( 1.37279) = $274.558